OpenStudy (anonymous):

if the parent function is y=1/x describe the change in the equation y=1/x+4

OpenStudy (ybarrap):

let f(x) = y = 1/x Then f(x) + 4 = y + 4 = 1/x + 4. Therefore, the change in f(x) in going from y to y + 4 is f(x) + 4. Graphically, the entire function, 1/x is shifted "upwards" by 4 units.

OpenStudy (anonymous):

thank you!

OpenStudy (ybarrap):

yw